OpenStudy (anonymous):

The measure of the angles of a triangle are in extended ratio 2:2:8. What is the measure of the smallest angle?

OpenStudy (bibby):

what do the interior angles of a triangle add up to?

OpenStudy (anonymous):

are the interior angles the 2s?

OpenStudy (bibby):

they're the 2, 2, and 8

OpenStudy (anonymous):

ohhhh ok ok i see, it would be 12

OpenStudy (bibby):



OpenStudy (bibby):

they're a ratio. it doesn't mean that the angles are 2, 2, and 8 degrees

OpenStudy (bibby):

it means that the measure of the smallest if 4 times smaller than the largest

OpenStudy (anonymous):

ok thats starting to make sense

OpenStudy (bibby):

so for example, let's do this with supplementary angles  the ratio of the larger angle to the smaller one is 8:1 together they're 9x so we can do 9x=180 x=20

OpenStudy (bibby):

and then plug it back into the ratio to figure out the angle measures. so 8*20=160 1*20=20 you can confirm they both add to 180

OpenStudy (bibby):

you have to do the same thing here with the triangle

OpenStudy (bibby):

\(2x+2x+8x=180\)

OpenStudy (anonymous):

Ok that makes sense

OpenStudy (anonymous):

thank you for your help!

OpenStudy (bibby):

anytime post your answer when you're here if you want me to double check